OpenStudy (anonymous):

does anyone know how to apply a transformation to a unit circle to get an ellipse?

OpenStudy (jamesj):

You want to dilate one of the axis. Suppose you dilate the x-axis. Then the transformation will be \[ \left( \begin{array}{cc} \lambda && 0 \\ 0 && 1 \end{array} \right) \] Under that transformation, the unit circle \[ x^2 + y^2 = 1 \] becomes \[ \lambda^2x^2 + y^2 = 1 \] which is an ellipse.

OpenStudy (anonymous):

the transformation i'm trying to apply is to get the ellipse x^2/36+y^2/12=1.. how might i go about doing that?

OpenStudy (jamesj):

well, under the transformation I wrote down, the ellipse can also be written as \[ \frac{x^2}{(1/\lambda)^2} + y^2 = 1 \] So what value of lambda would give you the first term in your ellipse? Now think about how to scale y to get your second term. What then does the matrix look like that scales both of those terms appropriately?

OpenStudy (anonymous):

soo.. it would look like (1/36 0 ) (0 1/12) ?

OpenStudy (jamesj):

almost. Notice we square x and y, and their results in the linear transformation.

OpenStudy (jamesj):

got it?

OpenStudy (anonymous):

ohh so 1/6 and 1/sqrt12?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

am i allowed to ask another question lol?

OpenStudy (jamesj):

Sure, post it on the left. Protocol now says you give me "Good answer" ;-)